# secReferences

### v1

@@ 3.2 Prediction of individual selection probability
Selection probability $\delta_{v}$ is a central quantity in defining both individual fairness and our new fairness. Since a KPD program is not static but dynamic, a natural statistical question to ask is how to predict individual selection probabilities before more incompatible donor-patient pairs enter the pool for the next round of exchange allocation. We provide a solution based on sample splitting. The following discussion can accommodate any fairness criteria within our general framework.
Our method utilizes historical data of incompatible donor-patient pairs independent of the current and future pairs. Suppose the historical pool consists of incompatible donor-patient pairs denoted as $\{\tilde{v}_{1},...,\tilde{v}_{N_{0}}\}$, and the current pool consists of pairs denoted as $\{v_{1},...,v_{N_{1}}\}$. Assume that the exchange allocation occurs when the size of the incompatible pairs pool reaches $N$, where $N_{1}<N<N_{0}+N_{1}$. The prediction procedure can be described in Algorithm 1 below.
@@
2:  Output: Prediction of $\delta_{v}$ for each $v\in\{v_{1},\dots,v_{N_{1}}\}$.
@@
6:     Based on the graph $(V^{(b)},E^{(b)})$, solve (4) to obtain the selection probability of $\delta_{v}^{(b)}$ for each $v\in\{v_{1},\dots,v_{N_{1}}\}$.
@@
8:  Return: mean and quantiles of $\{\delta_{v}^{(1)},...,\delta_{v}^{(B)}\}$ as the mean prediction and interval prediction of $\delta_{v}$, for each $v\in\{v_{1},\dots,v_{N_{1}}\}$.
In practice, the computational complexity could be very high to enumerate the cycles or relevant subsets of $\{v_{1},...,v_{N_{1}},\tilde{v}_{1}^{*},...,\tilde{v}_{N-N_{1}}^{*}\}$ for $B$, say 1000, times. When $M$ is small, it could be more computationally efficient to first enumerate the cycles or relevant subsets $\{v_{1},...,v_{N_{1}},\tilde{v}_{1}^{*},...,\tilde{v}_{N_{0}}^{*}\}$, and then filter the cycles or relevant subsets with vertices in $\{v_{1},...,v_{N_{1}},\tilde{v}_{1}^{*},...,\tilde{v}_{N-N_{1}}^{*}\}$ for each replication. However, when $N_{0}$ is extremely large, it could be impossible to enumerate the cycles or relevant subsets $\{v_{1},...,v_{N_{1}},\tilde{v}_{1}^{*},...,\tilde{v}_{N_{0}}^{*}\}$. One solution is to split $\{\tilde{v}_{1},...,\tilde{v}_{N_{0}}\}$ into disjoint subsets of appropriate size and to enumerate the cycles or relevant subsets within each subset accordingly.
Price of the new fairness criterion under random graph models
In this section, we establish theoretical guarantees for the price of fairness associated with our new fairness criterion using a random graph model that incorporates ABO blood types. We describe the model assumptions below.
@@
Recall that a donor and a patient are compatible if they match in both blood type and HLA. We assume blood type compatibility follows standard medical guidelines: AB patients can receive kidneys from donors of any blood type, A and B patients can receive from donors of their own type or type O, while O patients can only receive from type O donors. We assume HLA compatibility follows binomial distributions. Specifically, we randomly assign each patient a PRA score, representing the probability of being HLA incompatible with any donor. We assume the PRA scores can take discrete values $\{r_{1},\dots,r_{M}\}$, which also determine sensitization levels $\{r_{1},\dots,r_{M}\}$; and we assume HLA compatibility between different donor-patient pairs is independent. The vertex set $V$ is formed by independently drawing donor-patient pairs from an underlying population, keeping only incompatible pairs until a total of $N$ incompatible pairs is reached. The edge set $E$ is determined by the compatibility between donor-patient pairs in $V$. Regarding the utility assignment, we assume $u(c)$ only depends on the induced subgraph of vertices in $c$, independent of the protected feature of vertices in $c$.
@@ 4.1 Efficiency loss of subgroup prioritization
$\max_{x_{c}\in\{0,1\}}\quad\sum_{c\in\mathcal{C}}x_{c}u(c)\quad\text{subject to}\quad\sum_{c\in\mathcal{C}}x_{c}1_{v\in c}\leq 1,\ \forall\ v\in V$ (7)
but here, we allow for general utility $u$ and a set of cycles or relevant subsets $\mathcal{C}$ with some length limits. Let $\mathcal{P}$ denote the set of indices $(b_{1},b_{2},i,j)$, where the subgroup $V_{b_{1},b_{2},i,j}$ should be prioritized over $V_{b_{1},b_{2},1-i,j}$. We consider the optimization problem under the constraint of prioritizing these subgroups in $\mathcal{P}$:
@@
subject $\text{subject to}&\quad\sum_{c\in\mathcal{C}}x_{c}1_{v\in c}\leq 1,\ \forall\ v\in V,\\ &\left\{\sum_{v\in V_{b_{1},b_{2},i,j}}\sum_{c\in\mathcal{C}}x_{c}1_{v\in c}-|V_{b_{1},b_{2},i,j}|\right\}\cdot\left\{\sum_{v\in V_{b_{1},b_{2},1-i,j}}\sum_{c\in\mathcal{C}}x_{c}1_{v\in c}\right\}=0,\ \forall(b_{1},b_{2},i,j)\in\mathcal{P}.\end{aligned}$ (9)
The constraint (9) implies that no patients in $V_{b_{1},b_{2},1-i,j}$ can be matched or all patients in $V_{b_{1},b_{2},i,j}$ must be matched, if the subgroup $V_{b_{1},b_{2},i,j}$ is prioritized over $V_{b_{1},b_{2},1-i,j}$. The following Proposition 2 is the main result of this subsection.
@@
For a random graph with size $N$, the difference between the maximums achieved in the optimization problem (7) and the optimization problem (8)–(9) is $o(N)$, almost surely as $N\to\infty$.
Proposition 2 shows that we can prioritize patients with either $A=1$ or $A=0$, given their blood type and sensitization level, with ignorable relative efficiency loss when the random graph is large. The result is useful to derive upper bounds for the price of our new fairness, defined as the relative overall utility loss due to the fairness constraint in (4). We present these results in the next subsection.
4.2 Upper bounds for price of the new fairness
@@ 4.2.1 Optimizing some general utilities
First, we apply Proposition 2 to the scenario with general utilities, allowing for potential recourse strategies. Let $\mu_{b_{1},b_{2},r,a}$ denote the probability of sampling an incompatible donor-patient pair with donor blood type $b_{1}$, patient blood type $b_{2}$, patient sensitization level $r$, and patient sensitive group level $a$. We can obtain a crude upper bound for the price of fairness in Proposition 3 below.
@@ 4.2.2 Maximizing the number of transplants without recourse strategies
Specifically, let $\mu_{a}$ represent the frequency probability of the protected feature level $A=a$. Define $\mu_{O\mid a}$, $\mu_{A\mid a}$, $\mu_{B\mid a}$, and $\mu_{AB\mid a}$ as the frequency probabilities of blood types O, A, B, and AB, respectively, within the protected feature level $A=a$. Under these assumptions, there exists a constant $c$ such that $\mu_{b_{1},b_{2},r,a}=cr\mu_{a}\mu_{b_{1}\mid a}\mu_{b_{2}\mid a}$ for all $b_{1},b_{2}\in{O,A,B,AB}$, $r\in{r_{1},\dots,r_{M}}$, and $a\in{0,1}$.
@@
$\begin{split}T_{a}(r)=&\mu_{a}r(\mu_{O\mid a}+\mu_{AB\mid a}-\mu_{O\mid a}\mu_{AB\mid a}+\mu_{A\mid a}^{2}+\mu_{B\mid a}^{2})+2\mu_{a}\mu_{A\mid a}\mu_{B\mid a},\\ S_{a}(r)=&T_{a}(r)+\mu_{a}\{\mu_{O\mid a}(1-\mu_{O\mid a})+\mu_{A\mid a}\mu_{AB\mid a}+\mu_{B\mid a}\mu_{AB\mid a}\},\\ Q(r)=&T_{1}(r)+T_{0}(r)+\phi_{B,AB}+\phi_{O,AB}+\phi_{A,AB}+\phi_{A,O}+\phi_{AB,O}+\phi_{O,B},\\ \end{split}$
for $a=0,1$ and $r\in\{r_{1},...,r_{M}\}$, and
$\begin{split}R_{a}=&\min\left\{\phi_{B,AB},2\mu_{a}\mu_{B\mid a}\mu_{AB\mid a}-\phi_{B,AB}\right\}+\min\left\{\phi_{O,AB}+\phi_{A,AB},2\mu_{a}\mu_{A\mid a}\mu_{AB\mid a}-\phi_{O,AB}-\phi_{A,AB}\right\}\\ &+\min\left\{\phi_{A,O}+\phi_{AB,O},2\mu_{a}\mu_{O\mid a}\mu_{A\mid a}-\phi_{A,O}-\phi_{AB,O}\right\}+\min\left\{\phi_{O,B},2\mu_{a}\mu_{O\mid a}\mu_{B\mid a}-\phi_{O,B}\right\},\\ \end{split}$
@@
$\max_{r\in\{r_{1},...,r_{M}\}}\max\left\{\frac{S_{1}T_{0}-S_{0}T_{1}-S_{0}R_{1}}{S_{1}Q},\frac{S_{0}T_{1}-S_{1}T_{0}-S_{1}R_{0}}{S_{0}Q},0\right\}$
@@
The condition $1.5\overline{\mu}_{A}>\overline{\mu}_{O}>\overline{\mu}_{A}>\overline{\mu}_{B}>\overline{\mu}_{AB}$ gives a mild constraint on the blood types distribution of the whole population. The condition $\overline{\gamma}<0.4$ implies that most patients in the population are not highly-sensitized. Both assumptions are standard and appear in Ashlagi and Roth (2011) and Dickerson et al. (2014). As an application of Proposition 4, we present the following two hypothetical examples to illustrate the asymptotic upper bounds for the price of fairness are very small. The data come from the distribution of blood types in the United States as of 2021, according to the American Red Cross.
@@
As the first example, suppose there are two ethnicity groups, 80% white American and 20% African American, with $\mu_{O\mid 1}=0.45,\mu_{A\mid 1}=0.4,\mu_{B\mid 1}=0.11,\mu_{AB\mid 1}=0.04$ and $\mu_{O\mid 0}=0.51,\mu_{A\mid 0}=0.26,\mu_{B\mid 0}=0.19,\mu_{AB\mid 0}=0.04$. Moreover, there are three sensitization levels $\{0.05,0.45,0.9\}$. Then, with high probability, the price of fairness converges to 0 for any $0.40.05<\overline{\gamma}<$ 0.4. As the second example, suppose there are two ethnicity groups, 90% white American and 10% Asian American, with $\mu_{O\mid 1}=0.45,\mu_{A\mid 1}=0.4,\mu_{B\mid 1}=0.11,\mu_{AB\mid 1}=0.04$ and $\mu_{O\mid 0}=0.4,\mu_{A\mid 0}=0.275,\mu_{B\mid 0}=0.255,\mu_{AB\mid 0}=0.07$. Moreover, there are five sensitization levels $\{0.05,0.25,0.45,0.65,0.9\}$. Then, with high probability, the price of fairness is lower than 0.01 for any $0.090.05<\overline{\gamma}<$ 0.09, and converges to 0 for any $0.400.09\leq\overline{\gamma}<$ 0.40. The two examples show that the price of fairness can be no greater than $1\%$ for large graphs in real practice.

### v2

@@ 3.3 Prediction of individual selection probability
The selection probability $\pi_{v}$ is a central quantity in defining both individual fairness and our new fairness criterion. Since a KPD program is not static but dynamic, a natural statistical question to ask is how to predict individual selection probabilities before more incompatible donor-patient pairs enter the pool for the next round of exchange allocation. We provide a solution based on sample splitting. The following discussion can accommodate any fairness criteria within our general framework.
Our method utilizes historical data of incompatible donor-patient pairs independent of the current and future pairs. Suppose the historical pool consists of incompatible donor-patient pairs denoted as $\{\tilde{v}_{1},\cdots,\tilde{v}_{N_{0}}\}$, and the current pool consists of pairs denoted as $\{v_{1},\cdots,v_{N_{1}}\}$. Assume that the exchange allocation occurs when the size of the incompatible pairs pool reaches $N$, where $N_{1}<N<N_{0}+N_{1}$. The historical pool $\{\tilde{v}_{1},\cdots,\tilde{v}_{N_{0}}\}$ aggregates incompatible donor-patient pairs collected across potentially multiple past KPD rounds rather than a single pool. In each iteration $b$, we sample a subset of these historical vertices to simulate a feasible past pool of size $N-N_{1}$. The edge set $E^{(b)}$ is then determined over the combined vertex set $V^{(b)}=\{v_{1},\cdots,v_{N_{1}}\}\cup\tilde{V}^{(b)}$, treating unobserved edges between vertices that never coexisted in historical rounds as absent. To construct $E^{(b)}$, we recompute donor-patient compatibilities among all vertices in $V^{(b)}$ using the same ABO and HLA matching criteria as in the real KPD graph. Each edge $(v_{i},v_{j})$ is included in $E^{(b)}$ if donor $v_{i}$ is compatible with patient $v_{j}$. The prediction procedure can be described in Algorithm 1 below.
@@
2:Output: Prediction of $\pi_{v}$ for each $v\in\{v_{1},\dots,v_{N_{1}}\}$.
@@
6:  Based on the graph $(V^{(b)},E^{(b)})$, solve (4) to obtain the selection probability of $\pi_{v}^{(b)}$ for each $v\in\{v_{1},\dots,v_{N_{1}}\}$.
@@
8:Return: mean and quantiles of $\{\pi_{v}^{(1)},\cdots,\pi_{v}^{(B)}\}$ as the mean prediction and interval prediction of $\pi_{v}$, for each $v\in\{v_{1},\dots,v_{N_{1}}\}$.
In practice, the computational cost can be high if we directly enumerate all cycles or relevant subsets within the simulated pool $\{v_{1},\cdots,v_{N_{1}},\tilde{v}_{1}^{(b)},\cdots,\tilde{v}_{N-N_{1}}^{(b)}\}$ for each of the $B$ replications (e.g., $B=1000$). When the number of protected-feature groups $M$ is small, it is more efficient to first enumerate the feasible cycles or subsets once using the entire combined vertex set $\{v_{1},\cdots,v_{N_{1}},\tilde{v}_{1},\cdots,\tilde{v}_{N_{0}}\}$, and then, for each replication $b$, retain only those cycles whose vertices appear in the sampled subset $\{v_{1},\cdots,v_{N_{1}},\tilde{v}_{1}^{(b)},\cdots,\tilde{v}_{N-N_{1}}^{(b)}\}$. This strategy avoids repeated enumeration while preserving correctness. When $N_{0}$ is extremely large, however, full enumeration over all historical vertices may still be infeasible. In such cases, the historical pool $\{\tilde{v}_{1},\cdots,\tilde{v}_{N_{0}}\}$ can be partitioned into several disjoint subsets of manageable size, and the enumeration of feasible cycles or relevant subsets can be performed separately within each subset. The results are then combined across subsets as needed.
Price of our new fairness criterion under random graph models
The price of our new fairness criterion quantifies the efficiency loss incurred when enforcing fairness constraints compared to the unconstrained optimal allocation. Formally, it is defined as the relative reduction in the total utility (e.g., expected number of transplants or any general utility) achieved under our fairness-constrained optimization versus the utility achieved by the unconstrained solution. This concept follows the standard “price of fairness” framework in the literature (e.g., Bertsimas et al. 2011), here adapted specifically to our new fairness criterion. In this section, we establish theoretical guarantees for the price of our fairness criterion using a random graph model that incorporates ABO blood types. We describe the model assumptions below.
@@
Recall that a donor and a patient are compatible if they match in both blood type and HLA. We assume blood type compatibility follows standard medical guidelines: AB patients can receive kidneys from donors of any blood type, A and B patients can receive from donors of their own type or type O, while O patients can only receive from type O donors. We assume HLA compatibility follows binomial distributions. Specifically, we randomly assign each patient a PRA score, representing the probability of being HLA incompatible with any donor. We assume the PRA scores can take discrete values $\{r_{1},\dots,r_{M}\}$, which also determine sensitization levels $\{r_{1},\dots,r_{M}\}$; and we assume HLA compatibility between different donor-patient pairs is independent. The vertex set $V$ is formed by independently drawing donor-patient pairs from an underlying population, keeping only incompatible pairs until a total of $N$ incompatible pairs is reached. Let $\mu_{b_{1},b_{2},r,a}$ denote the probability of sampling an incompatible donor-patient pair with donor blood type $b_{1}$, patient blood type $b_{2}$, patient sensitization level $r$, and patient protected group level $a$. Here, the parameters $\{\mu_{b_{1},b_{2},r,a}\}$ are fixed but unknown. The edge set $E$ is determined by the compatibility between donor-patient pairs in $V$. Regarding the utility assignment, we assume $u(c)$ only depends on the induced subgraph of vertices in $c$, independent of the protected feature of vertices in $c$. We also assume $u(c)/|c|$ is bounded by a constant.
In our random graph model, the PRA levels $\{r_{1},\ldots,r_{M}\}$ are treated as fixed parameters representing different sensitization categories rather than random draws from a continuous distribution. Specifically, each vertex $v_{i}$ (representing a donor-patient pair) is assigned a PRA level $r_{i}$ from this finite set according to empirical proportions observed in real KPD data (e.g., Saidman et al. 2006). Hence, the PRA levels are not sampled uniformly over $[0,1]$ but follow a categorical distribution reflecting realistic sensitization patterns.
In our theoretical analysis, we adopt a simplified random graph model where each edge is independently present with a constant probability $p\in(0,1)$. This assumption is primarily technical and facilitates the use of classical random graph results (e.g., Erdős and Rényi 1968) to establish asymptotic matching properties. In real kidney exchange programs, however, the probability that an edge exists between two vertices depends on medical compatibility, particularly blood type and PRA sensitization level, so that the effective edge probabilities are heterogeneous across vertex pairs. Nevertheless, the assumption of a constant $p$ captures the key structural property required in our proofs: the existence of a sufficiently dense graph where each vertex has an expected number of compatible partners growing linearly in $N$. Under any model where the minimal edge probability remains bounded away from zero across vertex types (as is empirically the case in our simulations), the asymptotic results and efficiency bounds derived here continue to hold qualitatively. Therefore, while the assumption of a constant $p$ simplifies the theoretical exposition, it does not materially affect the validity of the conclusions regarding the asymptotic price of fairness.
@@ 4.1 Efficiency loss of subgroup prioritization
But here, we allow for a general utility $u$ and a set of cycles or relevant subsets $\mathcal{C}$ with some length limits. The term “general utility” refers to a flexible specification of the objective beyond the traditional “number of transplants” criterion. Specifically, $u(c)$ can represent any valid utility function that incorporates additional clinical or policy considerations, such as predicted graft survival, donor-recipient age matching, or sensitization adjustments, while still assigning a scalar utility value to each feasible exchange cycle $c$. Regardless of how $u(c)$ is defined, all utilities enter the optimization problem through binary decision variables, so the resulting formulation remains an integer programming problem.
Let $\mathcal{P}$ denote the set of indices $(b_{1},b_{2},i,j)$, where the subgroup $V_{b_{1},b_{2},i,j}$ should be prioritized over $V_{b_{1},b_{2},1-i,j}$. For fixed $\epsilon\geq 0$, we consider the optimization problem under the constraint of prioritizing these subgroups in $\mathcal{P}$:
@@
$\begin{aligned} \text{subject to}&\quad\sum_{c\in\mathcal{C}}x_{c}1_{v\in c}\leq 1,\ \forall\ v\in V,\\ &\sum_{v\in V_{b_{1},b_{2},i,j}}\sum_{c\in\mathcal{C}}x_{c}1_{v\in c}-|V_{b_{1},b_{2},i,j}|\leq\epsilon N\quad\mathrm{or}\quad\sum_{v\in V_{b_{1},b_{2},1-i,j}}\sum_{c\in\mathcal{C}}x_{c}1_{v\in c}\leq\epsilon N,\ \forall\ (b_{1},b_{2},i,j)\in\mathcal{P}.\end{aligned}$ (7)
The problems (5) and (6)–(7) share the same objective function. The constraint (7) means that at most $\epsilon N$ in $V_{b_{1},b_{2},1-i,j}$ can be matched or at most $\epsilon N$ patients in $V_{b_{1},b_{2},i,j}$ can be unmatched, if the subgroup $V_{b_{1},b_{2},i,j}$ is prioritized over $V_{b_{1},b_{2},1-i,j}$. We refer to (5) as the unconstrained problem, aside from the vertex-disjointness condition, and to (6)–(7) as the constrained problem.
The following Proposition 1 is the main result of this subsection.
@@
Consider the random graph model in Section 4, and fix any sequence $\varepsilon_{N}\downarrow 0$. For each $N$, let $\mathrm{Opt}(G_{N})$ be the optimum of the unconstrained problem (5), and let $\mathrm{Opt}^{\mathrm{prio}}_{\varepsilon_{N}}(G_{N})$ be the optimum of the constrained problem (6)–(7) with $\epsilon=\epsilon_{N}$. Then, with probability one as $N\to\infty$,
$\mathrm{Opt}(G_{N})-\mathrm{Opt}^{\mathrm{prio}}_{\varepsilon_{N}}(G_{N})=o(N).$
Proposition 1 shows that we can prioritize patients with either $A=1$ or $A=0$, given their blood type and sensitization level, with ignorable relative efficiency loss when the random graph is large. The result is useful to derive upper bounds for the price of our new fairness criterion, defined as the relative overall utility loss due to the fairness constraint in (4). We present these results in the next subsection.
4.2 Upper bounds for price of our new fairness criterion
@@ 4.2.1 Optimizing some general utilities
First, we apply Proposition 1 to the scenario with general utilities, allowing for potential recourse strategies. We can obtain a crude upper bound for the price of fairness in Proposition 2 below.
@@ 4.2.2 Maximizing the number of transplants without recourse strategies
Specifically, let $\mu_{a}$ represent the frequency probability of the protected feature level $A=a$. Define $\mu_{O\mid a}$, $\mu_{A\mid a}$, $\mu_{B\mid a}$, and $\mu_{AB\mid a}$ as the frequency probabilities of blood types O, A, B, and AB, respectively, within the protected feature level $A=a$. Under these assumptions, there exists a constant $c$ such that $\mu_{b_{1},b_{2},r,a}=cr\mu_{a}\mu_{b_{1}\mid a}\mu_{b_{2}\mid a}$ for all $b_{1},b_{2}\in\{O,A,B,AB\}$, $r\in\{r_{1},\dots,r_{M}\}$, and $a\in\{0,1\}$.
@@
[…]
for $a=0,1$ and $r\in\{r_{1},\cdots,r_{M}\}$, and
[…]
@@
$\max_{r\in\{r_{1},\cdots,r_{M}\}}\max\left\{\frac{S_{1}T_{0}-S_{0}T_{1}-S_{0}R_{1}}{S_{1}Q},\frac{S_{0}T_{1}-S_{1}T_{0}-S_{1}R_{0}}{S_{0}Q},0\right\}$
@@
Intuitively, the quantities in Proposition 3 summarize how much matching potential each protected group has, and how much of that potential can be reallocated when we impose our fairness constraint. For a fixed protected group level $A=a$ and sensitization level $R=r$, $T_{a}(r)$ is the expected fraction of incompatible pairs in group $a$ (at level $r$) that are fully matched in the optimal unconstrained allocation. It aggregates all blood-type combinations that are essentially “always used” in the optimal construction (e.g., X–X pairs, A–B/B–A pairs, and the core part of AB–O, A–O, B–O exchanges) and therefore contribute deterministically to the match count. $S_{a}(r)$ is the expected total fraction of incompatible pairs in group $a$ (at level $r$) that belong to blood-type classes that can potentially participate in exchanges under the Ashlagi–Roth optimal structure. It includes both the fully matched categories counted in $T_{a}(r)$ and the partially matched categories (such as B–AB, A–AB, O–A, O–B). Thus, the ratio $T_{a}(r)/S_{a}(r)$ represents the limiting selection probability for group $a$ at sensitization level $r$ in the unconstrained optimum.
At the population level, $Q(r)$ is the expected total fraction of incompatible pairs at sensitization level $r$ that are matched in the unconstrained optimal allocation, aggregated over both protected groups and all relevant blood-type combinations. It is the denominator in the expression for the asymptotic price of fairness, i.e., the benchmark efficiency against which we compare the fair allocation. Finally, $R_{a}$ captures the additional matching mass that can be shifted toward group $A=a$ within the partially matched blood-type classes while preserving feasibility. Each $\min\{\cdot,\cdot\}$ term in the definition of $R_{a}$ corresponds to a specific partially matched blood-type class (B–AB, A–AB, O–A/O–AB, O–B), and ensures we do not reassign more matches to group $a$ than either (i) the total number of matches available in that class, or (ii) the supply of compatible pairs in group $a$. In this sense, $R_{a}$ measures the “room to maneuver” for reallocating matches across protected groups without changing the blood-type structure of the exchange.
In combination, $T_{a}(r)$ and $S_{a}(r)$ describe the baseline matching rates for each group and sensitization level, $Q(r)$ encodes the total unconstrained efficiency, and $R_{a}$ quantifies how much we can adjust the allocation in favor of each group when enforcing our fairness constraint. The bound in Proposition 3 then compares the best fair allocation achievable under these structural limits with the unconstrained optimum, yielding an explicit asymptotic upper bound on the price of fairness.
The condition $1.5\overline{\mu}_{A}>\overline{\mu}_{O}>\overline{\mu}_{A}>\overline{\mu}_{B}>\overline{\mu}_{AB}$ imposes a mild and empirically realistic constraint on the population-level blood type distribution, while the condition $\overline{\gamma}<0.4$ implies that most patients are not highly sensitized. Both assumptions are standard in the kidney exchange literature and appear in Ashlagi and Roth (2011) and Dickerson et al. (2014). Moreover, these assumptions are consistent with real U.S. blood type frequencies reported by the American Red Cross (2021), approximately 45% type O, 40% type A, 11% type B, and 4% type AB, and with observed sensitization patterns in national KPD data, where the majority of patients exhibit moderate or low PRA levels.
@@
As an application of Proposition 3, we present two illustrative examples showing that the asymptotic upper bounds on the price of fairness are very small under empirically supported distributions. In the first example, suppose there are two ethnicity groups, 80% White American and 20% African American, with $\mu_{O\mid 1}=0.45$, $\mu_{A\mid 1}=0.40$, $\mu_{B\mid 1}=0.11$, $\mu_{AB\mid 1}=0.04$ and $\mu_{O\mid 0}=0.51$, $\mu_{A\mid 0}=0.26$, $\mu_{B\mid 0}=0.19$, $\mu_{AB\mid 0}=0.04$. Moreover, there are three sensitization levels $\{0.05,0.45,0.9\}$. Then, with high probability, the price of fairness converges to zero for any $0.05<\overline{\gamma}<0.4$.
In the second example, suppose there are two ethnicity groups, 90% White American and 10% Asian American, with $\mu_{O\mid 1}=0.45$, $\mu_{A\mid 1}=0.40$, $\mu_{B\mid 1}=0.11$, $\mu_{AB\mid 1}=0.04$ and $\mu_{O\mid 0}=0.40$, $\mu_{A\mid 0}=0.275$, $\mu_{B\mid 0}=0.255$, $\mu_{AB\mid 0}=0.07$. Moreover, there are five sensitization levels $\{0.05,0.25,0.45,0.65,0.9\}$. Then, with high probability, the price of fairness is less than $0.01$ for any $0.05<\overline{\gamma}<0.09$ and converges to zero for any $0.09\leq\overline{\gamma}<0.40$.
These two examples demonstrate that, under empirically grounded population characteristics, the price of fairness in large kidney exchange graphs is negligible, typically below 1%, confirming that fairness constraints can be imposed with minimal efficiency loss in real-world settings.